### v1

# IPAD: Inverse Prompt for AI Detection - A Robust and Explainable LLM-Generated Text Detector

Abstract
Large Language Models (LLMs) have attained human-level fluency in text generation, which complicates the distinguishing between human-written and LLM-generated texts. This increases the risk of misuse and highlights the need for reliable detectors. Yet, existing detectors exhibit poor robustness on out-of-distribution (OOD) data and attacked data, which is critical for real-world scenarios. Also, they struggle to provide explainable evidence to support their decisions, thus undermining the reliability. In light of these challenges, we propose  IPAD (Inverse Prompt for AI Detection), a novel framework consisting of a  Prompt Inverter that identifies predicted prompts that could have generated the input text, and a  Distinguisher that examines how well the input texts align with the predicted prompts. We develop and examine two versions of  Distinguishers. Empirical evaluations demonstrate that both  Distinguishers perform significantly better than the baseline methods, with version2 outperforming baselines by 9.73% on in-distribution data (F1-score) and 12.65% on OOD data (AUROC). Furthermore, a user study is conducted to illustrate that IPAD enhances the AI detection trustworthiness by allowing users to directly examine the decision-making evidence, which provides interpretable support for its state-of-the-art detection results.
IPAD: Inverse Prompt for AI Detection - A Robust and Explainable LLM-Generated Text Detector
Zheng Chen1thanks: Authors contributed equally to this work.,Yushi Feng211footnotemark: 1,Changyang He3,Yue Deng1,Hongxi Pu4,Bo Li1 1Hong Kong University of Science and Technology, Hong Kong 2The University of Hong Kong, Hong Kong 3Max Planck Institute, Germany 4University of Michigan, United States Correspondence: zchenin@connect.ust.hk
1 Introduction
Large Language Models (LLMs), characterized by their massive scale and extensive training data  Chen et al. (2024), have achieved significant advances in natural language processing (NLP)  Ouyang et al. (2022); Veselovsky et al. (2023); Wu et al. (2025). However, with the advanced capabilities of LLMs, they are subject to frequent misused in various domains, including academic fraud, the creation of deceptive material, and the generation of fabricated information Ji et al. (2023); Pagnoni et al. (2022); Mirsky et al. (2023), which underscores the critical need to distinguish between human-written text (HWT) and LLM-generated text (LGT) Pagnoni et al. (2022); Yu et al. (2025); Kirchenbauer et al. (2023).
However, due to their sophisticated functionality, LLMs pose significant challenges in the robustness of current AI detection systems Wu et al. (2025). The existing detection systems, including commercial ones, frequently misclassify texts as HWT Price and Sakellarios (2023); Walters (2023) and generate inconsistent results when analyzing the same text using different detectors Chaka (2023); Weber-Wulff et al. (2023). Studies show false positive rates reaching up to 50% and false negative rates as high as 100% in different tools Weber-Wulff et al. (2023) when dealing with out-of-distribution (OOD) datasets.
Another critical issue with the existing AI detection systems is their lack of verifiable evidence Halaweh and Refae (2024), as these tools typically provide only simple outputs like "likely written by AI" or percentage-based predictions Weber-Wulff et al. (2023). The lack of evidence prevents users from defending themselves against false accusations Chaka (2023) and hinders organizations from making judgments based solely on the detection results without convincing evidences Weber-Wulff et al. (2023). This problem is particularly troublesome not only because the low accuracy of such systems as mentioned before, but also due to the consequent inadequate response to LLM misuse, which can lead to significant societal harm Stokel-Walker and Van Noorden (2023); Porsdam Mann et al. (2023); Shevlane et al. (2023); Wu et al. (2025). These limitations highlight the pressing need for more reliable, explainable and robust detectors.
Figure 1: The overall workflow of our proposed IPAD framework
In this paper, we propose IPAD (Inverse Prompt for AI Detection), a novel framework comprising two key components as shown in Figure 1: a  Prompt Inverter that reconstructs prompts from input text, and a  Distinguisher that classifies text as HWT or LGT. We consider and examine two distinct approaches for the Distinguisher: the  Prompt-Text Consistency Verifier evaluates direct alignment between predicted prompts and input text, while the  Regeneration Comparator examines contents similarity by comparing input texts with the corresponding regenerated texts. Our framework introduces a paradigm shift in AI text detection by establishing an interpretable pipeline that reveals the underlying step-by-step reasoning process, therefore it enhances both detection robustness and explainability. Through comprehensive experiments comparing these two  Distinguishers, we demonstrate their respective strengths and limitations, providing new insights into how different text characteristics affect detection performance.
Empirical evaluations demonstrate that both  Distinguishers significantly surpass baseline methods, with the  Regeneration Comparator outperforming baselines by 9.73% (F1-score) on in-distribution data and 12.65% (AUROC) on out-of-distribution (OOD) data. Additionally, the  Regeneration Comparator exhibits better performance than the  Prompt-Text Consistency Verifier on attacked data with 3.78% (F1-score), and slightly better on OOD data with 0.13% (F1-score). Furthermore, a user study indicates that IPAD enhances the AI detection experience and trustworthiness by allowing users to directly examine its decision-making evidence, which includes the predicted prompts and regenerated texts, and hence provide transparent and interpretable support for its state-of-the-art detection results. The code will be available at https://github.com/Bellafc/IPAD-Inver-Prompt-for-AI-Detection-.
2 Methodology
In this section, we illustrate our method step by step. First, we introduce the overall workflow. After that, we demonstrate the details of supervised fine-tuning (SFT) the Prompt Inverter and Distinguisher .
2.1 Workflow
IPAD consists of a  Prompt Inverter and a  Distinguisher, both fine-tuned on Microsoft’s open model Phi3-medium-128k-instruct, which together form a complete detection workflow as illustrated in Figure 1. For the Distinguisher, we develop two models and examine them in Section  3.
The Input Text ($T$) is either human-written (HWT) or LLM-generated (LGT), and it is processed by the Prompt Inverter to predict the most likely prompt that could have generated it. This Predicted Prompt ($P$) is assumed to be the input that an LLM would have used to produce the text.
$P=f_{\text{inv}}(T)$
where $f_{\text{inv}}$ stands for Prompt Inverter.
For the next step, the Predicted Prompt ($P$) is fed into an LLM (we use ChatGPT, i.e. gpt-3.5-turbo by default, and other LLMs for evaluations), to generate a corresponding Regenerated Text ($T^{\prime}$).
$T^{\prime}=f_{\text{LLM}}(P)$
After that, we consider two Distinguishers. The first one is  Prompt-Text Consistency Verifier, in which the Input Text ($T$) and the Predicted Prompt ($P$) are passed to the model.
The  Prompt-Text Consistency Verifier determines whether the Predicted Prompt ($P$) can reasonably generate the given Input Text ($T$) using an LLM. The model outputs either a  "yes" or  "no" response. If the Predicted Prompt ($P$) is likely to produce the Input Text ($T$) when fed into the LLM, the model is expected to output  "yes", indicating that the Input Text ($T$) is likely LGT. Conversely, if the Predicted Prompt ($P$) does not align well with the Input Text ($T$), the model outputs  "no", suggesting that the Input Text ($T$) is less likely to have been generated by the LLM with the Predicted Prompt ($P$), and is therefore more likely to be HWT.
$S=f_{\text{PTCV}}(T,P)$
where $f_{\text{PTCV}}$ stands for Prompt-Text Consistency Verifier in the  Distinguisher.
The second Distinguisher is  Regeneration Comparator, which considers both the Input Text ($T$) and the Regenerated Text ($T^{\prime}$).
The  Regeneration Comparator determines whether the Input Text ($T$) aligns with the Regenerated Text ($T^{\prime}$), and then outputs either a "yes" or "no" response. If the Input Text ($T$) is LGT, the model is expected to output "yes," which indicates that both the Input Text ($T$) and the Regenerated Text ($T^{\prime}$) were generated by an LLM from similar prompts. Conversely, if the Input Text ($T$) is HWT, the model is expected to output "no," which signifies that the Input Text ($T$) is meaningfully distinct from the Regenerated Text ($T^{\prime}$) and thus unlikely to have been generated by an LLM.
$S=f_{\text{RC}}(T,T^{\prime})$
where $f_{\text{RC}}$ stands for  Regeneration Comparator in the Distinguisher.
Finally, for both Distinguishers,
$\hat{Y}=\begin{cases}\text{LGT},&\text{if }S=\text{Yes}\\ \text{HWT},&\text{if }S=\text{No}\end{cases}$
where $\hat{Y}$ is the final decision of the Input Text ($T$).
2.2 Datasets
2.2.1 Prompt Inverter
The datasets used to fine-tune the Prompt Inverter include several widely adopted resources in the field. These are:
  • Instructions-2M Morris et al. (2024), a collection of 2 million user and system prompts, from which we used 30,000 prompts.
  • ShareGPT Zhang et al. (2024b), an open platform where users share ChatGPT prompts and responses, from which we used 500 samples.
  • Unnatural Instructions Zhang et al. (2024b), a dataset of diverse, creative instructions generated by OpenAI’s text-davinci-002, from which we used 500 samples.
  • OUTFOX dataset Koike et al. (2024), which contains 15,400 essay problem statements, student-written essays, and LLM-generated essays.
The first three datasets aims to enhance the general querying capability of the Prompt Inverter, and are all released under the MIT license. All the samples we used are the same to the samples randomly selected in  Zhang et al. (2024a). The last dataset aims to enhance the familiarity of the Prompt Inverter with the data of the essay to detect the LLM-generated essays, and are created and examined by  Koike et al. (2024), We specifically used the LLM-generated essays and problem statements for this supervised fine-tuning (SFT).
For all 45,400 training pairs, the format is standardized as follows: Instruction: "Predict the prompt of the Input Text." Input: {LGT} or {HWT} Output: {Corresponding prompt}.
2.2.2 Distinguishers
Given that essay data are diverse, we utilize only the OUTFOX dataset Koike et al. (2024). To adapt this dataset for training our Distinguisher, we enhance it to align with the model’s requirements. The original dataset consists of 14,400 training triplets of essay problem statements, student-written essays, and LLM-generated essays. To further process the data, we apply the Prompt Inverter to both student-written and LLM-generated essays, generating corresponding Predicted Prompts. These Predicted Prompts are then used to regenerate texts via ChatGPT, i.e. gpt-3.5-turbo.
The final dataset is structured as follows:
Distinguisher version1 - Prompt-Text Consistency verifier: Instruction:"Can LLM generate text2 through the prompt text1? " Input: text 1: {Predicted Prompt}; text 2: {LGT} (or {HWT}) Output: yes (or no)
Distinguisher version2 - Regeneration Comparator: Instruction: "Text 1 is generated by an LLM. Determine whether Text 2 is also generated by an LLM with a similar prompt." Input: text 1: {Regenerated Text}; text 2: {LGT} (or {HWT}) Output: yes (or no)
Following this procedure, we construct a total of 28,800 training samples, with an equal distribution of positive and negative examples (14,400 each).
2.3 Training
The supervised fine-tuning (SFT)  Wei et al. (2022) process is performed on a dataset comprising the above-mentioned 45,400 pairs for Prompt Inverter and 28,800 pairs for both Distinguishers. We utilize Microsoft’s open model, phi3-medium-128k-instruct, and we use low-rank adaptation (LoRA) method  Hu et al. (2022) on the LLaMA-Factory framework [^1] https://huggingface.co/papers/2403.13372 Zheng et al. (2024). We train it using six A800 GPUs for 20 hours for Prompt Inverter, 7 hours for Distinguisher version1, and 4 hours for Distinguisher version2.
3 Experiments
We investigate the following questions through our experiments:
  • $\bullet$
    Assess the robustness of IPAD (using various LLMs as generators, comparing with other detectors, and evaluating on out-of-distribution (OOD) datasets).
  • $\bullet$
    Independently analyze the necessity and effectiveness of the Prompt Inverter and the Distinguishers.
  • $\bullet$
    Explore the explainability of IPAD (through a user study and analysis of linguistic differences between prompts generated by HWT and LGT).
3.1 Robustness of IPAD
3.1.1 Evaluation Baselines and Metrics
The in-distribution experiments refer to the testing results presented in  Koike et al. (2024), where the data aligns with the training data used for the IPAD  Distinguishers, thereby serving as our baseline. The OOD experiments refer to the DetectRL baseline  Wu et al. (2024), which is a comprehensive benchmark consisting of academic abstracts from the arXiv Archive (covering the years 2002 to 2017) [^2] http://kaggle.com/datasets/spsayakpaul/arxiv-paper-abstracts/data, news articles from the XSum dataset  Narayan et al. (2018), creative stories from Writing Prompts  Fan et al. (2018), and social reviews from Yelp Reviews Zhang et al. (2015). It also employs three attack methods to simulate complex real-world detection scenarios, which includes the prompt attacks, paraphrase attacks, and perturbation attacks Wu et al. (2024). All the testing sets have 1,000 samples in our experiments.
The  Area Under Receiver Operating Characteristic curve (AUROC) is widely used for assessing detection method  Mitchell et al. (2023) because it considers the True Positive Rate (TPR) and False Positive Rate (FPR) across different classification thresholds. Since our models predicts binary labels, we follow the  Wilcoxon-Mann-Whitney statistic Calders and Jaroszewicz (2007), and the formula is shown in appendix  A. The  AvgRec is the average of  HumanRec and  MachineRec. In our evaluation,  HumanRec is the recall for detecting Human-written texts, and  MachineRec is the recall for detecting LLM-generated texts Li et al. (2024). The  F1 Score provides a comprehensive evaluation of detector capabilities by balancing the model’s Precision and Recall. We use  AvgRec and  F1 on in-distribution data, and we use  AUROC for OOD data to align the test benchmarks for the same dataset.
3.1.2 Robustness across different LLMs
The results of IPAD for detecting the dataset OUTFOX Koike et al. (2024) across LLMs are presented in Table 1 and Table 2, respectively. They show that both versions are highly robust across various LLMs, while  Regeneration Comparator is a bit more efficient.
As for  Regeneration Comparator, when the original generator and re-generator are the same model, the performance is optimal. However, even when the re-generator is different from the original generator, the results remain impressive with ChatGPT used as the re-generator. These results imply that, in practical applications, it is possible to use a common set of LLMs as re-generators. If one or more correponding  Distinguishers from different LLMs classify the results as ’yes’, it can be inferred that the text is likely to be LGT, whereas if all  Distinguishers classify the results as ’no’, the text is more likely to be HWT. Furthermore, for applications aiming to save computational resources and improve efficiency, using ChatGPT as the sole re-generator still yields robust performance across all tested models.
Original Generator Metrics (%)
HumanRec MachineRec AvgRec F1
ChatGPT 98.00% 99.80% 98.90% 98.89%
GPT-3.5 97.20% 99.90% 98.55% 98.53%
Qwen-turbo 98.00% 98.10% 98.05% 98.05%
Llama-3-70B 98.00% 100.00% 99.00% 98.99%
Table 1: IPAD with  Prompt-Text Consistency Verifier performance on different LLMs
Original Generator Re-Generator Metrics (%)
HumanRec MachineRec AvgRec F1
ChatGPT ChatGPT 99.70% 100.00% 99.85% 99.85%
GPT-3.5 GPT-3.5 98.00% 100.00%  99.00%  99.00%
ChatGPT 97.00% 100.00% 98.50% 98.50%
Qwen-turbo Qwen-turbo 98.00% 98.40%  98.20%  98.20%
ChatGPT 99.70% 94.40% 97.05% 97.13%
Llama-3-70B Llama-3-70B 96.60% 100.00% 98.30% 98.30%
ChatGPT 99.70% 99.40%  99.55%  99.55%
Table 2: IPAD with  Regeneration Comparator performance on different LLMs
3.1.3 Comparison of IPAD with other detectors in and out of distribution
Table 3 compares the performance of two versions of IPAD with other detection methods in the OUTFOX dataset with and without attacks Koike et al. (2024). The results show that both versions of IPAD generally outperform other detectors, while that IPAD with  Prompt-Text Consistency Verifier for detecting ChatGPT with DIPPER attack performs worse. These results imply that IPAD with  Regeneration Comparator demonstrates superior robustness compared to alternative detection methods in the OUTFOX dataset with and without attacks.
Original Generator Detection Methods Metrics (%)
HumanRec MachineRec AvgRec F1
ChatGPT RoBERTa-base 93.80% 92.20% 93.00% 92.90%
RoBERTa-large 91.60% 90.00% 90.80% 90.70%
HC3 detector 79.20% 70.60% 74.90% 73.80%
OUTFOX 97.80% 92.40% 95.10% 95.00%
IPAD version1 98.00% 99.80% 98.90% 98.89%
IPAD version2 99.70% 100.00% 99.85% 99.85%
GPT-3.5 RoBERTa-base 93.80% 92.00% 92.90% 92.80%
RoBERTa-large 92.60% 92.00% 92.30% 92.30%
HC3 detector 79.20% 85.00% 82.10% 82.60%
OUTFOX 97.60% 96.20% 96.90% 96.90%
IPAD version1 97.20% 99.90% 98.55% 98.53%
IPAD version2 97.00% 100.00% 98.50% 98.50%
ChatGPT with DIPPER Attack RoBERTa-base 93.80% 89.20% 91.50% 91.30%
RoBERTa-large 91.60% 97.00% 94.30% 94.40%
HC3 detector 79.20% 3.40% 41.30% 5.50%
OUTFOX 98.60% 66.20% 82.40% 79.00%
IPAD version1 98.00% 75.10% 86.55% 87.93%
IPAD version2 99.70% 95.40% 97.55% 97.60%
ChatGPT with OUTFOX Attack RoBERTa-base 93.80% 69.20% 81.50% 78.90%
RoBERTa-large 91.60% 56.20% 73.90% 68.30%
HC3 detector 79.20% 0.40% 39.80% 0.70%
OUTFOX 98.80% 24.80% 61.80% 39.40%
IPAD version1 98.00% 95.40% 96.70% 96.74%
IPAD version2 99.70% 98.00% 98.85% 98.86%
Table 3: Comparison of IPAD with other detectors on in-distribution data, where  IPAD version1 stands for  IPAD with  Prompt-Text Consistency Verifier and  IPAD version2 stands for  IPAD with  Regeneration Comparator
Table 4 presents the performance of various detection methods on OOD datasets to assess their generalizability, where the baseline data refer to DetectRL  Wu et al. (2024). The results demonstrate that IPAD with  Regeneration Comparator consistently outperforms all other baselines in all OOD datasets with and without attacks. In contrast, IPAD with  Prompt-Text Consistency Verifier exhibits strong performance on OOD datasets without attacks but shows a noticeable drop in effectiveness when subjected to attacks. For instance, while it achieves competitive results on datasets like XSum (99.90%) and Writing (99.20%), its performance against attacks, such as Prompt Attack (86.90%) and Paraphrase Attack (82.72%), is significantly lower than IPAD with  Regeneration Comparator. This suggests that IPAD with  Regeneration Comparator demonstrates better generalizability and robustness.
OOD Datasets or attack type Detection Methods
LRR Fast-DetectGPT Rob-Base IPAD with version1 IPAD version2
Arxiv 48.17% 42.00% 81.06% 84.47% 98.60%
XSum 48.41% 45.72% 76.81% 99.90% 98.90%
Writing 58.70% 51.13% 86.29% 99.20% 95.80%
Review 58.21% 54.55% 87.84% 98.50% 89.30%
Avg. for non-attacked datasets 53.37% 48.35% 83.00% 95.52% 95.65%
Prompt Attack 54.97% 43.89% 92.81% 86.90% 93.05%
Paraphrase Attack 49.23% 41.15% 90.02% 82.72% 95.89%
Perturbation Attack 53.62% 44.38% 92.12% 94.96% 95.32%
Avg. for attacked datasets 52.61% 43.14% 91.65% 88.26% 94.75%
Avg. 53.04% 46.12% 86.70% 92.41% 95.26%
Table 4: The performance of IPAD in generalization assessment (AUROC). The selected detectors are evaluated on OOD data, all sourced from and processed using the DetectRL baseline, where  IPAD version1 stands for  IPAD with  Prompt-Text Consistency Verifier and  IPAD version2 stands for  IPAD with  Regeneration Comparator.
3.1.4 Robustness conclusion
Our experimental results demonstrate that both IPAD versions exhibit strong performance across different LLMs, outperforming existing detection methods and maintaining robustness on OOD datasets. The IPAD with  Regeneration Comparator outperforming baselines by 9.73% (F1-score) on in-distribution data and 12.65% (AUROC) OOD data. Notably, IPAD with  Regeneration Comparator achieves significantly better performance than IPAD with  Prompt-Text Consistency Verifier in attack scenarios of 3.78% (F1-score). While IPAD with  Prompt-Text Consistency Verifier performs robustly in standard settings, its performance declines when facing attacks. The calculation of these statistics are shown in Appendix  B.
3.2 Necessity and Effectiveness of Prompt Inverter and Distinguishers
3.2.1 Necissity of the Prompt Inverter and Distinguishers
To prove that it is necessary to fine-tune on IPAD with IPAD with  Prompt-Text Consistency Verifier and  Regeneration Comparator, we conducted ablation study to use the same finetune method on only  input texts and only  predicted prompts. The instructions are  "Is this text generated by LLM?", and  "Prompt Inverter predicts prompt that could have generated the input texts. Is this prompt predicted by an input texts written by LLM?", respectively.
The results shown in Figure  2 from the ablation study show that fine-tuning on either only the  input text or only the  predicted prompt leads to poor performance. This underscores the importance of fine-tuning on a combination of both the input text and predicted prompt, as explored in the  Prompt-Text Consistency Verifier, or on the input text and regenerated text, as examined in the  Regeneration Comparator, for more effective detection.
Figure 2: Ablation Study Results. The  IPAD version1 stands for  IPAD with  Prompt-Text Consistency Verifier and  IPAD version2 stands for  IPAD with  Regeneration Comparator.
3.2.2 The effectivenss of the IPAD Prompt Inverter
We use DPIC Yu et al. (2024) and PE Zhang et al. (2024c) as baseline methods for prompt extraction. DPIC employs a zero-shot approach using the prompt states in Appendix  C, while PE uses adversarial attacks to recover system prompts.
In our evaluation, we tested 1000 LGT and 1000 HWT samples. We use only in-distribution data for testing since only these datasets include original prompts. The metrics are all tested on comparing the similarity of the original prompts and the predicted prompts. The results shown in Table  5 illustrate that IPAD consistently outperforms both DPIC and PE across all four metrics (BartScore Yuan et al. (2021), Sentence-Bert Cosine Similarity Reimers and Gurevych (2019), BLEU Papineni et al. (2002), and ROUGE-1 Lin (2004)), which highlight the effectiveness of the IPAD  Prompt Inverter.
Evaluation Bart-large-cnn Sentence-Bert BLEU ROUGE-1
LGT
DPIC -2.12 0.46 5.61E-05 0.04
PE -2.23 0.58 3.21E-04 0.25
IPAD  -1.84  0.69  0.24  0.51
HWT
DPIC -2.47 0.42 8.75E-06 0.06
PE -2.39 0.53 2.56E-08 0.13
IPAD  -2.22  0.57  1.30E-01  0.39
Table 5: Comparison of the IPAD Prompt Inverter with other prompt extractors
3.2.3 The Effectiveness of the IPAD Distinguishers
To examine the effectiveness of the IPAD  Distinguishers, we conducted a comparison study using the same dataset but different distinguishing methods. The first and second methods employed Sentence-Bert  Reimers and Gurevych (2019) and Bart-large-cnn  Yuan et al. (2021) to compute the similarity score between the input texts and the regenerated texts. We selected thresholds that maximized AvgRec, which were 0.67 for Sentence-Bert and -2.52 for Bart-large-cnn. The classification rule is that the texts with scores greater than the threshold will be classified as LGT, while the texts with scores less than or equal to the threshold will be classified as HWT.
The third and fourth methods involved directly prompting ChatGPT as follows:
Instruction:  "Text 1 is generated by an LLM. Determine whether Text 2 is also generated by an LLM with a similar prompt. Answer with only YES or NO."  Input:  "Text 1: {Regenerated Text}; Text 2: {LGT} or {HWT}".
and  Instruction:  "Can LLM generate text2 through the prompt text1? Answer with only YES or NO." with  Input:  "Text 1: {Predicted Prompt}; Text 2: {Input text}".
The final results demonstrated that the other distinguishing methods performed worse than the two IPAD  Distinguishers, highlighting the superior effectiveness of the IPAD  Distinguishers.
Distinguish Method HumanRec MachineRec AvgRec F1
Sentence-Bert (Threshold 0.67) 61.20% 95.20% 78.20% 63.51%
Bart-large-cnn (Threshold -2.52) 42.60% 97.20% 69.90% 43.96%
Prompt to ChatGPT version 1 33.20% 64.50% 48.85% 44.77%
Prompt to ChatGPT version 2 12.50% 100% 56.25% 12.50%
IPAD version 1 98.00% 99.80% 98.90% 98.10%
IPAD version 2  99.70%  100%  99.85%  99.70%
Table 6: Comparison of Different Distinguishers, where  IPAD version1 stands for  IPAD with  Prompt-Text Consistency Verifier and  IPAD version2 stands for  IPAD with  Regeneration Comparator.
3.3 Explanability Assessment of IPAD
3.3.1 Different Linguistic Features of HWT prompts and LGT prompts
This subsection of the evaluation aims to explore the linguistic features of prompts generated by HWT and LGT through the Prompt Inverter. We analyzed 1000 samples generated by HWT and 1000 samples generated by LGT, which are randomy selected from both in-distribution data and OOD.
The analysis is first conducted using the Linguistic Feature Toolkik (lftk) [^3] https://lftk.readthedocs.io/en/latest/, a commonly used general-purpose tool for linguistic features extraction, which provides a total of 220 features for text analysis. Upon applying this toolkit, we identified 20 features with significant differences in average values between the two groups, out of which 3 features showed statistically significant differences with p-values less than 0.05. These 3 differences can be summarized as one main aspects:  syntactic complexity. Beyond these, we referred to the LIWC framework  [^4] https://www.liwc.app/, which defines 7 function words variables and 4 summary variables. By comparing the difference, two of these 11 features is significantly distinguishable:  the pronoun usage and  the level of analytical thinking.
One of the primary distinctions between the HWT prompts and the LGT prompts is sentence complexity. LGT prompts are typically more complex, characterized by longer sentence lengths (mean value of 1.514 and 1.794), higher syllable counts (mean values of total syllabus three are 1.572 and 3.042), and more stop-words (mean values of 9.88 and 10.045). HWT prompts, on the other hand, are characterized by shorter, less complex sentences that are easier to process and understand, as examples shown in Appendix D Figure 3.
Beyond the differences in syntactic complexity, we also explored variables in LIWC. We did the difference comparison by using HWT and LGT prompts as inputs for ChatGPT, for example, instructing with the prompts ’determine the pronoun usage of this sentence, answer first person, second person, or third person’ and ’determine the level of analytical thinking of these sentences, answer a number from 1 to 5’. The results show that there are distinguish difference in pronoun usage and analytical thinking level. The HWT prompts frequently use second-person pronouns (e.g., ’you’) - 75 occurrences per 1,000 prompts - due to the subjective tone often employed in HWT. In contrast, LGT prompts primarily feature first- and third-person pronouns, with second-person pronouns appearing only 2 per 1,000 prompts. LGT prompts typically present instructions and questions in a more objective manner. As shown in Appendix  D Figure  4, LGT prompts show higher analytical thinking levels than HWT prompts. With level 1 as the lowest and level 5 as the highest, LGT has 68.9% of level 4 and 24.3% of level 5, but HWT has only 48.0% of level 4, and 0.8% of level 5. It suggests that LGT prompts encourage more analytical thinking, while HWT prompts tend to focus more on concrete examples, with less emphasis on critical analysis, as examples shown in Appendix D Figure  5.
3.4 User Study
To assess the explainability improvement of IPAD, we designed an IRB-approved user study with ten participants evaluating one HWT and one LGT article. We used IPAD version 2 due to its superior OOD performance and attack resistance. Participants compared three online detection platforms with screenshots shown in Appendix E [^5] https://www.scribbr.com/ai-detector/ [^6] https://quillbot.com/ai-content-detector [^7] https://app.gptzero.me/ with IPAD’s process (which displayed input texts, predicted prompts, regenerated texts, and final judgments). After evaluation, users rated IPAD on four key explainability dimensions. Transparency received strong ratings (40%:5, 60%:4), with users appreciating the visibility of intermediate processes. Trust scores were more varied (10%:3, 70%:4, 20%:5), but IPAD was generally considered more convincing than single-score detectors. Satisfaction was mixed (30%:3, 30%:4, 40%:5), with users acknowledging better detection but raising concerns about energy efficiency since IPAD runs three LLMs. Debugging received unanimous 5s, as users could easily analyze the predicted prompt and regenerated text to verify the decision-making process. If needed, users could refine the generated content by adjusting instructions, such as specifying a word count, making IPAD a more effective and user-friendly tool compared to black-box detectors.
4 Related Work
4.1 AI detectors Methods and challenges
AI text detection methods can be broadly categorized into four approaches Wu et al. (2025): watermarking, statistics-based methods, neural-based methods, and human-assisted methods.
Watermarking technology inserts specific patterns into training datasets Shevlane et al. (2023); Gu et al. (2022) or manipulates the model output during inference to embed a watermark Lucas and Havens (2023). However, watermarking needs to access of the LLM deployment and can face attacks, such as identifying and erasing the watermark Hou et al. (2024).  Statistics-based methods analyze inherent textual features to identify language patterns Kalinichenko et al. (2003); Hamed and Wu (2023), but their effectiveness depends on corpus size and model diversity Wu et al. (2025). Some other statistical methods use n-gram probability divergence Yang et al. (2024b) or similarity between original and revised texts Mao et al. (2024); Zhu et al. (2023) while still face robustness challenges under adversarial attacks Wu et al. (2025)Neural-based methods such as RoBERTa Liu et al. (2020), Bert Devlin et al. (2019), and XLNet Yang et al. (2019) have been robust in domain-specific tasks. Adversarial learning techniques are increasingly being used Yang et al. (2024a) to increase effectiveness in attacked datasets.
In addition to automated methods, human involvement plays a key role in detecting AI-generated text Wu et al. (2025).  Human-assisted detection leverages human intuition and expertise to identify inconsistencies such as semantic errors and logical flaws that may not be easily caught by algorithms Uchendu et al. (2023); Dugan et al. (2023). Moreover, given the challenges of current AI detection tools, which often lack verifiable evidence Chaka (2023), human involvement becomes even more critical to ensure the reliable and explainable detection.
4.2 Prompt Inverter techniques and applications
Prompt extraction techniques aim to reverse-engineer the prompts that generate specific outputs from LLMs. Approaches include black-box methods like output2prompt Zhang et al. (2024a), which extracts prompts based on model outputs without access to internal data, and logit-based methods like logit2prompt Mitka (2024), which rely on next-token probabilities but are constrained by access to logits. Adversarial methods can bypass some defenses but are model-specific and fragile Zhang et al. (2024d). Despite the success of some zero-shot LLM-inversion based methods Li and Klabjan (2024); Yu et al. (2024), they are mostly naive usage of prompting LLMs, which makes them poor in prompt extraction accuracy and robustness.
5 Conclusion
This paper introduces  IPAD (Inverse Prompt for AI Detection), a framework consisting of a  Prompt Inverter that identifies predicted prompts that could have generated the input text, and a  Distinguisher that examines how well the input texts align with the predicted prompts. This design enables explainable evidence chains tracing unavailable in existing black-box detectors. Empirical results show that IPAD surpasses the baselines on all in-distribution, OOD, and attacked data. Furthermore, the  Distinguisher (version2) -  Regeneration Comparator outperforms the  Distinguisher (version1) -  Prompt-Text Consistency Verifier, especially on OOD and attacked data. While the local alignment in veresion1 approach provides explicit interpretability, it is more sensitive to adversarial attacks. In contrast, the global distribution in veresion2 matching approach implicitly learns generative LLM’s distributional properties, which offers more robustness while maintaining explainability. This insight suggests that combining self-consistency checks of generative models with multi-step reasoning for evidential explainability holds promise for future AI detection systems in real-world scenarios. A user study reveals that IPAD enhances trust and transparency by allowing users to examine decision-making evidence. Overall, IPAD establishes a new paradigm for more robust, reliable, and interpretable AI detection systems to combat the misuse of LLMs.
6 Limitations
While IPAD demonstrates SOTA performance, two limitations warrant discussion: (1) The Prompt Inverter may not fully reconstruct prompts containing explicit in-context learning examples (e.g., formatted demonstrations), as it prioritizes semantic alignment over precise syntactic replication. (2) Since IPAD achieves satisfactory OOD performance (12.65% improvement over baselines) by only adopting essay writing datasets for the fine-tuning of Distinguishers, we strategically deferred the exploration of more datasets. We will incorporate a wider and more diverse range of data in future works to explore if it can enhance robustness even further, including: creative/news domains, and triplet data formats (i.e.,  "Can this {predicted prompt} generate the {Input text} using an LLM? One example generated by the predicted prompt is: {regenerated text}")
Acknowledgments
Appendix A AUROC formula
Since our model predicts binary labels, we follow the  Wilcoxon-Mann-Whitney statistic Calders and Jaroszewicz (2007) to calculate the Area Under Receiver Operating Characteristtic curve (AUROC):
$\text{AUC}(f)=\frac{\sum_{t_{0}\in\mathcal{D}^{0}}\sum_{t_{1}\in\mathcal{D}^{1}}\mathbf{1}[f(t_{0})<f(t_{1})]}{|\mathcal{D}^{0}|\cdot|\mathcal{D}^{1}|}$
where $\mathbf{1}[f(t_{0})<f(t_{1})]$ denotes an indicator function which returns 1 if $f(t_{0})<f(t_{1})$ and 0 otherwise. $\mathcal{D}^{0}$ is the set of negative examples, and $\mathcal{D}^{1}$ is the set of positive examples.
Appendix B Calculation of Summary Statistics
  • $\bullet$
    IPAD with  Regeneration Comparator outperforms the baselines by 9.73% on in-distribution data. As shown in Table 3, RoBERTa-base has the best average F1 score of (92.9% + 92.8% + 91.3% + 78.9%) / 4. In comparison, the average F1 score for IPAD version 2 is (99.85% + 98.5% + 97.6% + 98.86%) / 4, showing an improvement of 9.73%.
  • $\bullet$
    IPAD with  Regeneration Comparator outperforms the baselines by 12.65% on in-distribution data. As shown in Table 4, RoBERTa-base achieves the highest average AUROC score, but since the F1-score is not available for the baseline, we use the AUROC difference to calculate the improvement, which is (95.65% - 83%) = 12.65%.
  • $\bullet$
    IPAD with  Regeneration Comparator outperforms IPAD with  Prompt-Text Consistency Verifier by 0.13% on out-of-distribution (OOD) data. As shown in Table 4, IPAD version 2 has the highest AUROC of 95.65%, while IPAD version 1 has an AUROC of 95.52%, resulting in a 0.13% difference.
  • $\bullet$
    IPAD with  Regeneration Comparator outperforms IPAD with  Prompt-Text Consistency Verifier by 3.78% on attacked data. As shown in Table 3 (rows 3-4) and Table 4 (rows 6-8), IPAD version 2 achieves the best F1 score and AUROC scores. To calculate the overall attacked dataset score, we calculate the F1 scores for Table 4: 94.82%, 95.35%, 95.31% for IPAD version 2, and 83.58%, 88.34%, and 94.70% for IPAD version 1. The average F1 score difference is thus (94.82% + 95.35% + 95.31% - 83.58% - 88.34% - 94.70% + 97.60% + 98.86% - 97.55% - 98.85%) / 5 = 3.78%.
Appendix C DPIC (decouple prompt and intrinsic characteristics) Prompt Extraction Zero-shot Prompts
"I want you to play the role of the questioner. I will type an answer in English, and you will ask me a question based on the answer in the same language. Don’t write any explanations or other text, just give me the question. <TEXT>.".
Appendix D Linguistic Difference Examples
Figure  3 shows examples where HWT and LGT prompts with different sentence complexity. Figure  4 shows the results of analytical thinking level statistics. Figure  5 shows examples of using different personas and different analytical thinking levels.
Figure 3: Sentence Complexity Examples, where  HWT Prompt stands for prompt generated by the Prompt Inverter from HWT, and  LGT Prompt stands for prompt generated by the Prompt Inverter from LGT. The HWT Prompts have longer sentence lengths, more words with more than three syllabus (as shown in bold), and more stop-words (as shown with underline).
Figure 4: Comparison of different analytical thinking levels, with LGT has higher percentage of level 4 and level 5.
Figure 5: Examples that use different persona usage (above), and different analytical thinking levels (below left has level 2, and below right has level 5, they are prompts generated by the same problem statements).
Appendix E User Study
Figure  6 7 and  8 shows the screenshots of online AI detectors. Figure  9 shows the questionnaire questions. Figure  10 shows the user guide.
E.1 Online AI Detectors Screenshots
Figure 6: GPTZero Online Detector Screenshot
Figure 7: Quillbot Online Detector Screenshot
Figure 8: Scribbr Online Detector Screenshot
E.2 Questionnaire questions
Figure 9: Questionnaire questions
Figure 10: User Study User guide